How many grams of Aluminum chloride are produced when 734 mg of barium chloride react with an excess of Aluminum sulfate? show your work.
1
Expert's answer
2012-03-01T10:54:38-0500
0.734 & & & & & & & & & & & & & & m
3BaCl2+Al2(SO4)3=2AlCl3 +
3BaSO4
3*208
2*133.5

m=0.734*2*133.5/3*208=0.314 g